# Electromagnetic response and emergent topological orders in transition metal dichalcogenide MoTe2 bilayers

### v1

@@ I Introduction
The moiré band structure of t-MoTe2 and its topological properties are highly sensitive to the twist angle. For $2.0^{\circ}\lesssim\theta\lesssim 2.7^{\circ}$, the first few hole-doped moiré bands constructed from one valley all have the same Chern number $C=1$, while those constructed from the other valley have $C=-1$ instead, as required by time-reversal symmetry [7, 8, 9]. As the twist angle increases, several topological transitions occur. In the range $2.8^{\circ}\lesssim\theta\lesssim 3.9^{\circ}$, the topmost pair of bands per valley exhibit Chern numbers $C=\pm 1$, resembling a pair of Haldane bands [10].
@@
Motivated by this phenomenological picture, we carry out a comprehensive analysis of composite fermion states emerging from the Hofstadter spectrum induced by a mean-field uniform Chern-Simons flux in both Phases I and II. In this framework, the Hofstadter spectrum at finite external magnetic field can be directly transposed to the composite fermion picture, revealing a rich sequence of valley-polarized fractional Chern insulator (FCI) states. We map out the resulting incompressible phases as a function of filling fraction per valley and extract the fractional Hall conductivity from the topological properties (Chern numbers) of the filled composite fermion bands. A key novelty, in contrast to the conventional fractional quantum Hall effect (FQHE) in continuum Landau levels, is the emergence of composite fermion bands within a fractal Hofstadter spectrum, possessing Chern numbers distinct from those of Landau levels. These give rise to fractal FCIs [39] where the Hall conductivity and filling fraction deviate from the standard Jain hierarchy [40, 41] due to the significant influence of the moiré superlattice. Although these fractal FCIs tend to exhibit smaller composite fermion gaps, their experimental realization would mark an important advance in the study of fractionalized topological phases.
@@ II.1 The continuum model
$\begin{split}V_{b/t}^{(2)}(\mathbf{r})&=V_{2}\left[e^{i\left(\mathbf{g}_{1}+\mathbf{g}_{2}\right)\cdot\mathbf{r}}+e^{i\left(\mathbf{g}_{3}+\mathbf{g}_{4}\right)\cdot\mathbf{r}}+e^{i\left(\mathbf{g}_{5}+\mathbf{g}_{6}\right)\cdot\mathbf{r}}\right]\,,\\ T^{(2)}(\mathbf{r})&=w_{2}\left[e^{i\mathbf{g}_{1}\cdot\mathbf{r}}+e^{i\mathbf{g}_{4}\cdot\mathbf{r}}+e^{i\left(\mathbf{g}_{3}+\mathbf{g}_{2}\right)\cdot\mathbf{r}}\right]\,,\end{split}$ (3)
@@
The parameters in Eqs. (2) and (3) can be extracted by fitting to density functional theory (DFT) bands [7, 4]. In this work, we adopt the values ($a_{0}$, $m^{*}$, $V$, $\psi$, $w$, $V_{2}$, $w_{2}$) = ($0.3472\mathrm{~{}nm}$[53, 54], $0.62\mathrm{~{}m}_{\mathrm{e}}$, $20.51\mathrm{meV}$, $-61.49^{\circ}$,$-7.01\mathrm{meV}$, $-9.08\mathrm{meV}$, $11.08\mathrm{meV}$) at $\theta=2.1^{\circ}$ and ($0.355\mathrm{~{}nm}$[5], $0.62\mathrm{~{}m}_{\mathrm{e}}$, $9.45\mathrm{meV}$, $-85.23^{\circ}$,$-12.2\mathrm{meV}$, $24.99\mathrm{meV}$, $13.12\mathrm{meV}$) at $\theta=3.89^{\circ}$ from Ref.[6]. With these parameters, one can diagonalize the Hamiltonian in a plane wave basis, leading to the moiré band structures as shown in Fig.1(a) and (b).
@@ III.1 Coupling to magnetic fields: Hofstadter bands
$[R_{1},R_{2}]=-i\ell_{B}^{2}(\hat{{\bm{g}}}_{1}\times\hat{{\bm{g}}}_{2})\cdot\hat{z}=-i\frac{\sqrt{3}}{2}\ell_{B}^{2}\,,$ (7)
@@
Thus, upon expressing the guiding center coordinate as $y=y_{0}+(u\,p+j)\Delta$, where $0\leq y_{0}\leq\Delta$, $u\in\mathbb{Z}$ and $j=0,1,...,p-1$, (with $j\equiv j~{}\textrm{mod}$(p)) the action of the guiding center part $\mathcal{B}_{m_{1},m_{2}}$ in Eq.(6) reads
@@
$\begin{split}&\,\left\langle n_{2},y_{2}\right|e^{\pm i{\bm{q}}\cdot{\bm{\eta}}}\left|n_{1},y_{1}\right\rangle\\ &\,=\delta_{y_{1},y_{2}}\,\sqrt{\frac{n_{2}!}{n_{1}!}}e^{-x/2}(\frac{\pm iq_{-}\ell_{B}}{\sqrt{2}})^{n_{1}-n_{2}}L_{n_{2}}^{n_{1}-n_{2}}(x)\,,\end{split}$ (13)
for $n_{1}\geq n_{2}$, where $q_{-}=q_{x}-iq_{y}$, $x=|q_{-}|^{2}\ell_{B}^{2}/2$ and $L_{n_{2}}^{n_{1}-n_{2}}(x)$ is the Laguerre polynomial [See App.].
@@ III.2 Moiré Hofstadter bands at zero displacement field
The moiré Hofstadter bands at $u_{d}=0$ for Phases I and II are shown in Figs. 2(a) and Figs. 2(b). These features extend the analysis of Ref. [37, 38] by incorporating second harmonic contributions and the effects of displacement fields. These spectra illustrate the fractal band structure that emerges near the $K$ valley as a function of the normalized magnetic flux per moiré unit cell, $\phi/\phi_{0}$, for opposite orientations of the magnetic field applied perpendicular to the bilayer plane, where positive and negative flux correspond to magnetic fields along the $-z$ and $+z$ directions, respectively. We note that spectrum at the $K^{\prime}$ valley is obtained by time-reversal operation, which flips the chirality of the electronic states and the direction of the magnetic field. A comment of the Zeeman effect is in order before discussing the spectrum. For $\theta=2.1^{\circ}$, the magnetic field strength needed to produce $\phi_{0}$ is approximately $25T$. The g-factor for bilayer MoTe2 varies from 4.73(at 2K) to 2.54(at 70K)[56]. Thus, the Zeeman splitting at $\phi\approx\phi_{0}$ is estimated to be around $7$ meV, which is comparable to the first moiré band gap. Thus, with a magnetic field of this magnitude, first moiré band will be valley-polarized, in which case it is legitimate to consider only one valley.
The evolution of the bands exhibits a characteristic dependence on the magnetic field orientation, particularly in the behavior of the gap between the two topmost moiré bands. Specifically, the gap decreases for positive flux and increases for negative flux. This asymmetry correlates with the interplay between the external magnetic field direction and the average Berry curvature. In Phase I, the average Berry curvature of the top bands is positive, and the gap evolution closely mirrors that of conventional Landau levels. A similar trend is observed in Phase II, even though the top two bands exhibit opposite average Berry curvatures.
@@ III.3 Moiré Hofstadter bands at non-zero displacement field
We first note the displacement field can alter the topological character of the moiré bands at $\phi=0$. As shown in Fig. 3 and Fig. 4, increasing $u_{d}$ affects the bandwidths of the moiré bands leading to a sequence of phase transitions at $\phi=0$. In particular, the gap between the first and second bands closes at $u_{d}\approx 17.0$ meV ($D/\varepsilon_{0}\approx 242.9$ mV/nm) for $\theta\approx 2.1^{\circ}$, and at $u_{d}\approx 19.5$ meV ($D/\varepsilon_{0}\approx 278.6$ mV/nm) for $\theta\approx 3.89^{\circ}$. These estimates assume an inter-layer distance $d\approx 0.7$ nm and a relative permittivity $\varepsilon_{r}\approx 10$. After the gap closing, the Chern number of the first moiré band becomes zero.
@@ IV Composite fermion states: platforms for FCIs and FQSH
As discussed above, we attach uniform $b$ to $K$-valley Hamiltonian Eq. (4), and $-b$ to $K^{\prime}$-valley Hamiltonian. Given a particular field parametrized as $\phi=bA_{uc}=\frac{p}{q}\phi_{0}$ with $p$ and $q$ coprime integers (same as in Sec.III), the moiré hole filling fraction per valley is given by
@@
The composite fermion picture also allows for a new analysis on how the external $E$ field affects the FCI phases in t-MoTe2. To this end, we follow the recipe outlined in Sec.III.3 and investigate the composite fermion spectrum in the presence of $E$ field. We restrict to values of the field for which the first moiré band remains topological, which corresponds to $170<u_{d}<$ 17 meV for $\theta=2.1^{\circ}$ and $19.50<u_{d}<$ 19.5 meV for $\theta=3.89^{\circ}$. The results for $\theta=2.1^{\circ}$ and $\theta=3.89^{\circ}$ are presented in Fig. 7 and Fig. 8 respectively. In both cases, we find a common feature that the FCI gaps tend to decrease as the electric field increases.
@@
While the Jain states remain gapped with $u_{d}$ up to $17$meV, remarkably, we identified an electric field induced quantum phase transition (QPT) in the fractal FCI states. At $\nu=\frac{4}{5}$ and $\theta=2.1^{\circ}$, such QPT occurs twice as $u_{d}$ increases from 00 to around $17$meV, as demonstrated in Fig.7(a). The first transition between FCIs with $\sigma_{xy}=(3/5)(e^{2}/h)$ and $\sigma_{xy}=(2/5)(e^{2}/h)$ occurs at $u_{d}\approx 4$ meV, and a second transition back to $\sigma_{xy}=(3/5)(e^{2}/h)$ takes place at $u_{d}\approx 14$ meV. For $\nu=\frac{4}{5}$ and $\theta=3.89^{\circ}$, such a QPT only occurs once when $u_{d}$ sweeps from 00 to $19.5$meV, as shown in Fig.8(a).
@@ V Discussion and Outlook
We have also shown that the analysis of the fractal spectrum in the presence of magnetic field is of particular significance in characterizing the composite fermion states describing a host of FCI states in t-MoTe2. In particular, we have done a comprehensive analysis of the Hofstadter spectrum of composite fermions associated with a uniform Chern-Simons field that attaches two flux quanta per particle. We have identified both Jain FCI states at filling fractions $\nu=\nu_{\text{Jain}}=\frac{1}{3},\frac{2}{5},...,\frac{3}{5},\frac{2}{3}$, and a series of fractal FCI states at filling fractions $\nu=\nu_{\text{Fractal}}=\frac{1}{5},\frac{2}{9},\frac{4}{5},...$. Although the composite fermion bands for the Jain states are similar to the Landau levels with Chern numbers $C=\pm 1$, those for the fractal FCI states are more dispersive, with Chern numbers $C=2$ or $3$, which do not resemble Landau levels. Our analysis reveals that Jain states generally exhibit larger composite fermion gaps, indicating greater stability. While we have primarily focused on filling fractions $\nu<1$, where prominent composite fermion gaps are observed, an example for $21<\nu<$ 2 at $\theta\approx 2.1^{\circ}$ is discussed in Appendix D.
@@
Our discussion of Hofstadter spectrum of t-MoTe2 also serves as a departing point for future studies on the interplay between electron correlation and external magnetic field in this platform. A compelling question is to investigate how interacting symmetry breaking electronic orders, nested in the Hofstadter system, evolve and compete under large magnetic fluxes. Another open question involves the treatment of the Chern-Simons statistical flux field $b(\bm{r})$ in a self-consistent way. We have argued that for a robust FCI state, approximating $b(\bm{r})$ as some uniform $\bar{b}$ does not change the essential physics. In this situation, electron interaction stablizes the FCI state and the composite fermions are treated as non-interacting particles. However, this uniform field ansatz is biased towards liquid-like FCI states and thus may underestimate the role played by other electronic orders competing with the FCI state (such as CDW, or zero field versions of Hall-like crystals [66, 59]), for which a nonuniform $b(\bm{r})$ and nonzero residual interactions become important for understanding the competing ground state. Understanding these non-uniform flux-charge binding states is an outstanding question that warrants future investigation.
@@ Appendix A Hamiltonian in the Landau Level basis
Following [55], we perform flux attachment to electrons characterized by $\bm{\Pi}=\mathbf{p}-(-e)\mathbf{A}$ in an external magnetic field $\mathbf{B}=\nabla\times\mathbf{A}=-B\hat{\mathbf{z}}$.
@@
$\begin{split}\mathcal{B}_{1,0}|n,y\rangle&=e^{i\bm{g}_{1}\bm{R}}|n,y\rangle=e^{igR_{1}}|n,y\rangle\\ &=|n,y_{0}+(up+j-1)\Delta\rangle\,,\\ \mathcal{B}_{0,1}|n,y\rangle&=e^{i\bm{g}_{2}\bm{R}}|n,y\rangle=e^{igR_{2}}|n,y\rangle\\ &=e^{igy_{0}}e^{i2\pi\frac{q}{p}j}|n,y_{0}+(up+j)\Delta\rangle\,,\end{split}$ (31)
$\begin{split}\mathcal{B}_{-1,1}|n,y\rangle&=e^{i\bm{g}_{3}\bm{R}}|n,y\rangle=e^{-igR_{1}+igR_{2}}|n,y\rangle\\ &=e^{igy_{0}}e^{i2\pi\frac{q}{p}(j+1/2)}|n,y_{0}+(up+j+1)\Delta\rangle\,,\\ \mathcal{B}_{0,-1}|n,y\rangle&=e^{i\bm{g}_{5}\bm{R}}|n,y\rangle=e^{-igR_{2}}|n,y\rangle\\ &=e^{-igy_{0}}e^{-i2\pi\frac{q}{p}j}|n,y_{0}+(up+j)\Delta\rangle\,,\end{split}$ (32)
@@
$\begin{split}&V_{m\bm{g}_{1}+n\bm{g}_{2}}|n_{1},y\rangle\\ &=e^{i(m\bm{g}_{1}+n\bm{g}_{2})\cdot\bm{r}}|n_{1},y\rangle=\mathcal{A}_{m,n}\mathcal{B}_{m,n}|n_{1},y\rangle\\ &=e^{ingy_{0}}e^{i\frac{2\pi q}{p}n(j-\frac{m}{2})}\mathcal{A}_{m,n}|n_{1},y-m\Delta\rangle\\ &=e^{ingy_{0}}e^{i\frac{2\pi q}{p}n(j-\frac{m}{2})}\sum_{n_{2}}\langle n_{2},y-m\Delta|\mathcal{A}_{m,n}|n_{1},y-m\Delta\rangle\\ &\hskip 142.26378pt\times|n_{2},y-m\Delta\rangle\\ &=e^{ingy_{0}}e^{i\frac{2\pi q}{p}n(j-\frac{m}{2})}\sum_{n_{2}}F_{n_{2},n_{1}}(m\bm{g}_{1}+n\bm{g}_{2})|n_{2},y-m\Delta\rangle\,,\end{split}$ (34)
@@
$\begin{split}&\frac{\hbar^{2}}{m^{*}}\left(\Pi_{x}k_{x}^{t/b}+\Pi_{y}k_{y}^{t/b}\right)\\ =&\frac{\hbar^{2}}{m^{*}}\left(\frac{\eta_{y}}{\ell^{2}}k_{x}^{t/b}-\frac{\eta_{x}}{\ell^{2}}k_{y}^{t/b}\right)\\ =&\frac{\hbar^{2}}{m^{*}}\left(-i\frac{a-a^{\dagger}}{\sqrt{2}\ell}k_{x}^{t/b}-\frac{a+a^{\dagger}}{\sqrt{2}\ell}k_{y}^{t/b}\right)\\ =&\frac{\hbar^{2}}{m^{*}}\frac{1}{\sqrt{2}\ell}\left((-k_{y}^{t/b}+ik_{x}^{t/b})a^{\dagger}+(-k_{y}^{t/b}-ik_{x}^{t/b})a\right)\,,\end{split}$ (36)

### v2

@@ I Introduction
The moiré band structure of t-MoTe2 and its topological properties are highly sensitive to the twist angle. For $2.0^{\circ}\lesssim\theta\lesssim 2.7^{\circ}$, the first few hole-doped moiré bands constructed from one valley all have the same Chern number $C=1$, while those constructed from the other valley have C=1C=-1 instead, as required by time-reversal symmetry [7, 8, 9]. As the twist angle increases, several topological transitions occur. In the range $2.8^{\circ}\lesssim\theta\lesssim 3.9^{\circ}$, the topmost pair of bands per valley exhibit Chern numbers $C=\pm 1$, resembling a pair of Haldane bands [10].
@@
Motivated by this phenomenological picture, we carry out a comprehensive analysis of composite fermion states emerging from the Hofstadter spectrum induced by a mean-field uniform Chern-Simons flux in both Phases I and II. In this framework, the Hofstadter spectrum at finite external magnetic field can be directly transposed to the composite fermion picture, revealing a rich sequence of valley-polarized fractional Chern insulator (FCI) states. We map out the resulting incompressible phases as a function of filling fraction per valley and extract the fractional Hall conductivity from the topological properties (Chern numbers) of the filled composite fermion bands. A key novelty, in contrast to the conventional fractional quantum Hall effect (FQHE) in continuum Landau levels, is the emergence of composite fermion bands within a fractal Hofstadter spectrum, possessing Chern numbers distinct from those of Landau levels. These give rise to fractal FCIs [39] where the Hall conductivity and filling fraction deviate from the standard Jain hierarchy [40, 41] due to the significant influence of the moiré superlattice. It is worth noting that such emergent topological orders have also been numerically reported in [42] recently. Although these fractal FCIs tend to exhibit smaller composite fermion gaps, their experimental realization would mark an important advance in the study of fractionalized topological phases.
@@ II.1 The continuum model
[…]
@@
The parameters in Eqs. (2) and (3) can be extracted by fitting to density functional theory (DFT) bands [7, 4]. In this work, we adopt the values ($a_{0}$, $m^{*}$, $V$, $\psi$, $w$, $V_{2}$, $w_{2}$) = ($0.3472\mathrm{~nm}$[54, 55], $0.62\mathrm{~m}_{\mathrm{e}}$, $20.51\mathrm{meV}$, $61.49^{\circ}$$7.01\mathrm{meV}$, $9.08\mathrm{meV}$, $11.08\mathrm{meV}$) at $\theta=2.1^{\circ}$ and ($0.355\mathrm{~nm}$[5], $0.62\mathrm{~m}_{\mathrm{e}}$, $9.45\mathrm{meV}$, $85.23^{\circ}$$12.2\mathrm{meV}$, $24.99\mathrm{meV}$, $13.12\mathrm{meV}$) at $\theta=3.89^{\circ}$ from Ref.[6]. With these parameters, one can diagonalize the Hamiltonian in a plane wave basis, leading to the moiré band structures as shown in Fig.1(a) and (b).
@@ III.1 Coupling to magnetic fields: Hofstadter bands
$R_{1},R_{2}]=-i\ell_{B}^{2}(\hat{{\bm{g}}}_{1}\times\hat{{\bm{g}}}_{2})\cdot\hat{z}=-i\frac{\sqrt{3}}{2}\ell_{B}^{2}\,$, (7)
@@
Thus, upon expressing the guiding center coordinate as $y=y_{0}+(u\,p+j)\Delta$, where $0\leq y_{0}\leq\Delta$, $u\in\mathbb{Z}$ and j=0,1,,p1j=0,1,...,p-1, (with $j\equiv j~\textrm{mod}$(p)) the action of the guiding center part $\mathcal{B}_{m_{1},m_{2}}$ in Eq.(6) reads
@@
[…]
for $n_{1}\geq n_{2}$, where $q_{-}=q_{x}-iq_{y}$, $x=|q_{-}|^{2}\ell_{B}^{2}/2$ and $L_{n_{2}}^{n_{1}-n_{2}}(x)$ is the Laguerre polynomial (See Appendix A).
@@ III.2 Moiré Hofstadter bands at zero displacement field
The moiré Hofstadter bands at $u_{d}=0$ for Phases I and II are shown in Figs. 2(a) and Figs. 2(b). These features extend the analysis of Ref. [37, 38] by incorporating second harmonic contributions and the effects of displacement fields. These spectra illustrate the fractal band structure that emerges near the $K$ valley as a function of the normalized magnetic flux per moiré unit cell, $\phi/\phi_{0}$, for opposite orientations of the magnetic field applied perpendicular to the bilayer plane, where positive and negative flux correspond to magnetic fields along the z-z and +z+z directions, respectively. We note that spectrum at the $K^{\prime}$ valley is obtained by time-reversal operation, which flips the chirality of the electronic states and the direction of the magnetic field. A comment of the Zeeman effect is in order before discussing the spectrum. For $\theta=2.1^{\circ}$, the magnetic field strength needed to produce $\phi_{0}$ is approximately $25T$. The g-factor for bilayer MoTe2 varies from 4.73(at 2K) to 2.54(at 70K)[57]. Thus, the Zeeman splitting at $\phi\approx\phi_{0}$ is estimated to be around $7$ meV, which is comparable to the first moiré band gap. Thus, with a magnetic field of this magnitude, first moiré band will be valley-polarized, in which case it is legitimate to consider only one valley.
The evolution of the bands exhibits a characteristic dependence on the magnetic field orientation, particularly in the behavior of the gap between the two topmost moiré bands. Specifically, the gap decreases for positive flux and increases for negative flux. It’s worth noting that, in our convention, a positive flux corresponds to a magnetic field pointing in the negative $\hat{\mathbf{z}}-$direction. This asymmetry correlates with the interplay between the external magnetic field direction and the average Berry curvature. In Phase I, the average Berry curvature of the top bands is positive, and the gap evolution closely mirrors that of conventional Landau levels. A similar trend is observed in Phase II, even though the top two bands exhibit opposite average Berry curvatures.
@@ III.3 Moiré Hofstadter bands at non-zero displacement field
We first note the displacement field can alter the topological character of the moiré bands at $\phi=0$. As shown in Fig. 3 and Fig. 4, increasing $u_{d}$ affects the bandwidths of the moiré bands leading to a sequence of phase transitions at $\phi=0$. In particular, the gap between the first and second bands closes at $u_{d}\approx 17.0$ meV ($D/\varepsilon_{0}\approx{u_{d}/(e\cdot d)\cdot\varepsilon_{r}}=242.9$ mV/nm) for $\theta\approx 2.1^{\circ}$, and at $u_{d}\approx 19.5$ meV ($D/\varepsilon_{0}\approx 278.6$ mV/nm) for $\theta\approx 3.89^{\circ}$. These estimates assume an inter-layer distance $d\approx 0.7$ nm[58, 59, 60] and a relative permittivity $\varepsilon_{r}\approx 10$[61, 62]. After the gap closing, the Chern number of the first moiré band becomes zero.
@@ IV Composite fermion states: platforms for FCIs and FQSH
As discussed above, we attach uniform $b$ to $K$-valley Hamiltonian Eq. (4), and b-b to $K^{\prime}$-valley Hamiltonian. Given a particular field parametrized as $\phi=bA_{uc}=\frac{p}{q}\phi_{0}$ with $p$ and $q$ coprime integers (same as in Sec.III), the moiré hole filling fraction per valley is given by
@@
The composite fermion picture also allows for a new analysis on how the external $E$ field affects the FCI phases in t-MoTe2. To this end, we follow the recipe outlined in Sec.III.3 and investigate the composite fermion spectrum in the presence of $E$ field. We restrict to values of the field for which the first moiré band remains topological, which corresponds to $0<u_{d}<17$ meV for $\theta=2.1^{\circ}$ and $0<u_{d}<19.5$ meV for $\theta=3.89^{\circ}$. The results for $\theta=2.1^{\circ}$ and $\theta=3.89^{\circ}$ are presented in Fig. 7 and Fig. 8 respectively. In both cases, we find a common feature that the FCI gaps tend to decrease as the electric field increases.
@@
While the Jain states remain gapped with $u_{d}$ up to 1717meV, remarkably, we identified an electric field induced quantum phase transition (QPT) in the fractal FCI states. At $\nu=\frac{4}{5}$ and $\theta=2.1^{\circ}$, such QPT occurs twice as $u_{d}$ increases from 0 to around 1717meV, as demonstrated in Fig.7(a). The first transition between FCIs with $\sigma_{xy}=(3/5)(e^{2}/h)$ and $\sigma_{xy}=(2/5)(e^{2}/h)$ occurs at $u_{d}\approx 4$ meV, and a second transition back to $\sigma_{xy}=(3/5)(e^{2}/h)$ takes place at $u_{d}\approx 14$ meV. For $\nu=\frac{4}{5}$ and $\theta=3.89^{\circ}$, such a QPT only occurs once when $u_{d}$ sweeps from 0 to 19.519.5meV, as shown in Fig.8(a).
@@ V Discussion and Outlook
We have also shown that the analysis of the fractal spectrum in the presence of magnetic field is of particular significance in characterizing the composite fermion states describing a host of FCI states in t-MoTe2. In particular, we have done a comprehensive analysis of the Hofstadter spectrum of composite fermions associated with a uniform Chern-Simons field that attaches two flux quanta per particle. We have identified both Jain FCI states at filling fractions $\nu=\nu_{\text{Jain}}=\frac{1}{3},\frac{2}{5},...,\frac{3}{5},\frac{2}{3}$, and a series of fractal FCI states at filling fractions $\nu=\nu_{\text{Fractal}}=\frac{1}{5},\frac{2}{9},\frac{4}{5},$.... Although the composite fermion bands for the Jain states are similar to the Landau levels with Chern numbers $C=\pm 1$, those for the fractal FCI states are more dispersive, with Chern numbers $C=2$ or $3$, which do not resemble Landau levels. Our analysis reveals that Jain states generally exhibit larger composite fermion gaps, indicating greater stability. While we have primarily focused on filling fractions $\nu<1$, where prominent composite fermion gaps are observed, an example for $1<\nu<2$ at $\theta\approx 2.1^{\circ}$ is discussed in Appendix D.
@@
Besides the displacement field, temperature plays a critical role in observing the Hofstadter spectrum and the stability of the associated correlated phases. In moiré systems under external magnetic fields, both single-particle and fractionalized states associated with the fractal spectrum have been experimentally observed at temperatures ranging from 10 mK to 4.5 K[26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 71]. On the other hand, in the absence of external field, recent experiments[12, 18, 15] on FCIs in t-MoTe2 report thermal activation gaps of $\Delta_{\nu=2/3}=23$ K and $\Delta_{\nu=3/5}=15$ K at $\theta=3.7^{\circ}$[12], and $\Delta_{\nu=2/3}=$ 27 K at $\theta=3.64^{\circ}$[18]. For comparison, our composite fermion analysis at $\theta=3.89^{\circ}$ yields band gaps of 1.3 meV for $\nu=2/3$ and 2.1 meV for $\nu=3/5$, corresponding to thermal energy gaps of approximately $\Delta_{\nu=2/3}\approx 15$ K and $\Delta_{\nu=3/5}\approx 24$ K respectively, which are consistent with the experimentally reported scales.
Our discussion of Hofstadter spectrum of t-MoTe2 also serves as a departing point for future studies on the interplay between electron correlation and external magnetic field in this platform. A compelling question is to investigate how interacting symmetry breaking electronic orders, nested in the Hofstadter system, evolve and compete under large magnetic fluxes. Another open question involves the treatment of the Chern-Simons statistical flux field $b(\bm{r})$ in a self-consistent way. We have argued that for a robust FCI state, approximating $b(\bm{r})$ as some uniform $\bar{b}$ does not change the essential physics. In this situation, electron interaction stabilizes the FCI state and the composite fermions are treated as non-interacting particles. However, this uniform field ansatz is biased towards liquid-like FCI states and thus may underestimate the role played by other electronic orders competing with the FCI state[72, 65]. This is particularly relevant for the $\nu=1/3$ state, which emerges as one of the prominent gaps in our composite fermion analysis, but has not yet been observed experimentally. Several experimental studies on moiré systems, including t-MoTe2, suggest that the observed 1/3 state is likely a trivial correlated insulator[17, 73, 74]. This discrepancy may reflect the assumption inherent in our uniform flux attachment approximation, which favors fractional states over competing topologically trivial orders. A self-consistent treatment of the Chern-Simons flux attachment, allowing for spatial variations in density, could provide further insight into the competition between FCIs and CDWs in this regime. Understanding these non-uniform flux-charge binding states is an outstanding question that warrants future investigation.
@@ Appendix A Hamiltonian in the Landau Level basis
Following [56], we perform flux attachment to electrons characterized by $\bm{\Pi}=\mathbf{p}+e\mathbf{A}$ in an external magnetic field $\mathbf{B}=\nabla\times\mathbf{A}=-B\hat{\mathbf{z}}$.
@@
[…]
@@
[…]
@@
[…]